40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog  Not logged in ELOG logo
Entry  Mon Jan 4 10:46:01 2021, aaron, Lab Infrastructure, Temperature Sensing, fresh temperature sensor Screenshot_from_2021-01-04_16-13-53.png
    Reply  Tue Jan 5 16:43:44 2021, aaron, Lab Infrastructure, Temperature Sensing, fresh temperature sensor Screen_Shot_2021-01-05_at_16.46.14.png
Message ID: 2605     Entry time: Mon Jan 4 10:46:01 2021     Reply to this: 2606
Author: aaron 
Type: Lab Infrastructure 
Category: Temperature Sensing 
Subject: fresh temperature sensor 

Entered Mon Jan 4 10:45:56 2021 .

Temperature in the lab is 89 F at the sensor by the electronics rack. I had been planning to put our AD590s on a board today, unnecessary equipment is already off but I also turned off cominaux and the Sorensen DC supplies. There is hot air coming from HVAC.

I spent some time gathering components and checking out AD's notes on the sensor. I'm using the same components as Andrew/Johannes' old circuit, and putting it on one of Anchal's "generic op amp / diff amp" boards (D1900129-v2). After checking out the old circuit, I realized we still have it in the cryo lab -- it's just missing the AD590. I put the AD590 on a breadboard, made a few connectors, and confirmed that I get a sensible reading (83 K near the center of the room, a couple degrees hotter in/by the electronics rack where the sensor is located).

I noticed that the AC connection on one of our Tenma power supplies is loose (wobbling the cable switches the supply off/on). It's a dual supply, but since I wouldn't want to use it elsewhere I'm powering the AD590 with it (doesn't need to move, out of the way above the rack).

Here's the temperature (second trend) over 20 min, will post the overnight minute trends tomorrow.

Picked up the mess, changed the dust mats, exit

Attachment 1: Screenshot_from_2021-01-04_16-13-53.png  37 kB  Uploaded Mon Jan 4 16:14:20 2021  | Hide | Hide all
ELOG V3.1.3-